# Bose-Einstein condensation of THz photons in an optical microcavity with Landau-quantized electrons

### v1

@@ III.1 Condensate mode
Figure 6: Diagrams for photon scattering from the condensate to a satellite mode with $\omega_{\rm satell}=\omega_{c}$ and nonzero $\bm{q}_{\rm satell}$.
$\nu_{\rm satell}{=}\frac{2\pi}{\hbar}\!\!\!\sum_{\begin{subarray}{c}\bm{p}\\ \bm{q}_{\rm satell}\neq 0\\ {s=\pm 1}\end{subarray}}\!\!\!\left|\!\frac{\lambda_{n,n\shortminus s}^{-}(\bm{q}_{\rm satell})\lambda^{+}_{n\shortminus s,n}(\bm{0})}{E^{s}_{n}{+}s\hbar\omega_{c}{-}E_{n\shortminus s}^{-s}}\right|^{2}$
×$\times N^{\rm el}_{n,s}\left(1{-}N^{\rm el}_{n,s}\right)\delta(\hbar\omega_{\rm satell}{-}\hbar\omega_{c}).$ (12)
Here $\bm{q}_{\rm satell}$ denotes the 3D momentum of satellite mode photon.
III.2 Pumping modes
Besides being coupled to the condensate mode, pumping modes are exposed to the black body radiation source, which imposes the thermal occupation $N^{\rm ph}_{\rm therm}(T_{\rm pump})$ at temperature $T_{\rm pump}>T$:
$N^{\rm ph}_{\rm therm}(T_{\rm pump})=\left[{\exp{\left(\frac{\hbar\omega_{\rm pump}}{T_{\rm pump}}\right)}-1}\right]^{-1}.$
When considering the rate equation for the occupation of the pumping modes, we use the following equation:
$\frac{\partial_{t}N_{\rm pump}^{\rm ph}}{S}{=}W\!\left(\!1{-}\frac{N^{\rm ph}_{\rm pump}}{N^{\rm ph}_{\rm therm}}\!\right)\!{-}n_{0}(N^{\rm ph}_{\rm pump}\nu_{\rm pump}{-}\nu_{\rm reverse}).$ (13)
Here the condensate density $n_{0}=N_{0}/S$ is introduced with $S$ being the resonator area. In this equation we treat all the pumping modes collectively as a single mode assuming the pumping range broadening due to the magnetic field gradient to be narrow enough for the black body photon occupation not to change significantly over this interval.
In addition to the photon number conserving term $-n_{0}N_{\rm pump}\nu_{\rm pump}$ (see (13) ), we add a phenomenological term, which describes the finite flux to the pumping modes from the black body source. We use the following model expression for the coupling rate:
$W{=}\frac{\Delta N}{\Delta S\Delta t}$ ${=}\frac{c}{\sqrt{\varepsilon}}\!\int\!\frac{d^{3}p}{(2\pi\hbar)^{3}}\frac{1}{\exp(\hbar\omega_{\rm pump}/T_{\text{pump}})-1}.$ (14)
@@ III.3 Satellite modes
As mentioned above, when describing condensate coupling to satellite modes, the coupling rate does not depend on occupations of these modes. That is why we do not consider a kinetic equation for them.
@@ III.4 Stationary state
$\begin{cases}\partial_{t}n_{0}=n_{0}\left(N_{\rm pump}^{\rm ph}\nu_{\rm pump}-\nu_{\rm total}-\nu_{\rm mirror}\right),\\ \partial_{t}N_{\rm pump}^{\rm ph}/S=W\left(1-{N_{\rm pump}^{\rm ph}}/{N^{\rm ph}_{\rm therm}}\right)-n_{0}\left(N_{\rm pump}^{\rm ph}\nu_{\rm pump}-\nu_{\rm reverse}\right).\end{cases}$ (15)
We derive kinetic rates assuming that the condensate is represented by a single mode. However, hereinafter we consider $n_{0}$ to be a collective variable for the total occupation of the modes with low in-plane momentum (pictorially represented in Fig. 3 by a circle).
We use $\nu_{\rm total}=\nu_{\rm reverse}{+}\nu_{\rm satell}{+}\nu_{\rm loss}$ for the total loss rate of the condensate mode due to interaction with other ones. In addition, we introduced $\nu_{\rm mirror}(Q)=\omega_{c}/(2\pi Q)$ to describe photon outflow from the cavity with $Q$ being the Q-factor.
For the reasonable operation regime of the proposed device, one may safely consider the rates $\nu_{\rm pump},\;\nu_{\rm reverse},\;\nu_{\rm loss}$ and $\nu_{\rm satell}$ to be constant (independent of $n_{0}$), as proven in Appendix E.
Thus, the stationary solution of the system (15), which describes photonic condensate with macroscopic occupation of the corresponding mode is given as follows:
$N_{\rm pump}^{\rm ph}=\frac{\nu_{\rm total}+\nu_{\rm mirror}}{\nu_{\rm pump}},$ (16)
$n_{0}=\frac{W}{\nu_{\rm satell}+\nu_{\rm loss}+\nu_{\rm mirror}}\left(1-\frac{\nu_{\rm total}+\nu_{\rm mirror}}{\nu_{\rm pump}N^{\rm ph}_{\rm therm}}\right).$ (17)
@@ IV.1.1 Electron mobility
$\Gamma=\frac{e\hbar}{\mu_{e}m^{*}}.$ (18)
@@ IV.1.2 Designing a cavity
To create the desired mode configuration, as described in Sec. II, we propose using a heterostructure with high electron mobility inside a Bragg-reflector. This heterostructure hosts the layers of 2DEG (using multiple layers is helpful for increasing overall electron density). A possible pattern of the dielectric constant modulation is depicted in Fig. 7 (a) (half of the structure is shown with a symmetric counterpart implied to the left). In the same plot we show the typical light intensity profile into the properly designed cavity.
@@
For the sake of simplicity, the transmittance profile on panel (b) is approximated by a step function + a delta peak for further calculations. We use the following parameters: $\varepsilon=15.2$ (InSb), $\varepsilon_{a}=2.1$, $\varepsilon_{b}=5.9$, the values for TiO2/SiO2 Bragg reflector. We consider cavities with $L_{\rm ph}\in[4,40]\,\mu$m, which corresponds to $\omega_{c}\in[1,10]$ THz.
@@ IV.2 Operating regimes
The Fig. 9 shows the regions of condensate existence for various temperatures of the thermal emitter. Note that these regions (in blue) are bounded by both the absorption peaks and the line of threshold pumping level (in red), which is given as a result of solving the equation (see (15))
$N_{\rm pump}^{\rm ph}\nu_{\rm pump}(\omega_{c},H)-\nu_{\rm total}(\omega_{c},H)=0$ (21)
One may clearly see how the area of the condensate phase shrinks with decreasing temperature (note the reversed axis). The critical temperature is $T_{\rm pump}^{\rm crit}=1300$ K.
@@ IV.3 Device performance
Now we study the device performance in terms of emission power and efficiency. Since the useful THz radiation is the one that escapes the microcavity through the mirror, the stationary emission intensity $I_{c}=\hbar\omega_{c}\cdot\partial_{t}n_{0}\big{|}_{\rm mirror}$ is as follows:
$I_{c}=\frac{W\hbar\omega_{c}\nu_{\rm mirror}}{\nu_{\rm satell}+\nu_{\rm loss}+\nu_{\rm mirror}}\left(1-\frac{\nu_{\rm total}+\nu_{\rm mirror}}{\nu_{\rm pump}N^{\rm ph}_{\rm therm}}\right)$ (22)
@@
$I_{\rm in}\simeq\hbar\omega_{\rm pump}W\left(1-\frac{\nu_{\rm total}+\nu_{\rm mirror}}{\nu_{\rm pump}N^{\rm ph}_{\rm therm}}\right).$ (23)
@@
$\eta=\frac{I_{c}}{I_{\rm in}}\simeq\frac{\omega_{c}}{\omega_{\rm pump}}\frac{\nu_{\rm mirror}}{\nu_{\rm satell}+\nu_{\rm loss}+\nu_{\rm mirror}}.$ (24)
@@
$Q_{\rm min}=\frac{\omega_{c}}{2\pi(\nu_{\rm pump}N^{\rm ph}_{\rm therm}-\nu_{\rm satell}-\nu_{\rm reverse})}$ (25)
@@
The threshold Q-factor is quite high compared to the ones achieved in experiments ($\sim 10^{3}$)[26]. However, as described in the same Ref. [26], adding several more layers to the Bragg reflectors may increase $Q$ by orders of magnitude. In addition, one is not constrained by thermal pumping, which we use for demonstrative calculations. Non-equilibrium pumping sources may be used, i.e. photonic crystal structures that increase the flux on the desired pumping frequency. With increased pumping flux, the required Q-factor will be lowered.
Operation regime of the device is given by thin stripes in the figure. These are exactly the valleys between the peaks of condensate photon absorption by the 2DEG as depicted in Fig. 8. These stripes widen with the increase of magnetic field.
@@
Among the actions to take to increase the thermal input to the cavity is using THz (metallic) mirrors which reflect the light from wrong directions, forwarding it to the cavity.
In addition, when creating a device, one would like to decrease the energy, consumed by the lamp to produce radiation at higher frequencies. This may be achieved by placing the lamp inside a cavity made of high-frequency (IR frequency, visible light) Bragg mirrors, which trap the redundant radiation. Thus, the power consumption of the lamp drops after equilibrating with that cavity.
@@ V Conclusions
In this paper we presented a new mechanism for observing microcavity photonic BEC in THz range. We studied a system of microcavity photons exposed to pumping by a thermal light source. The temperature difference between the cavity and the pumping source instigates photonic flux to the former. Due to interphotonic interaction via Landau-quantized electrons in magnetic field, the system settles down in a stationary state with macroscopic occupation of a single transverse mode of the cavity.
The demonstrated possibility to populate a single transverse mode is necessary though insufficient condition for achieving full coherence. In contrast with cold-atom condensate, there is no direct spontaneous symmetry breaking because of negligible interaction between condensate photons. The condensate gain saturation mechanism is purely kinetic (limited influx from the pumping modes), which is closer to the condensation scenario in exciton-polariton condensate.
In this study we treated a bunch of low-lying modes of the condensate collectively. This definitely prevents the essential discussion on the thermalization dynamics and statistics of the condensate. That is why the estimation of the coherence degree of the photonic condensate and investigation of its thermalization dynamics is beyond the scope of the current paper. A detailed study of this issue (in the fashion of [27] for the photonic condensate in a dye-filled cavity) is a subject of future investigations.
@@ A.1 Electron-phonon
[…]
@@
$\psi^{y}_{m}(y,p_{x})$ $=\frac{H_{m}(\eta)e^{-\eta^{2}/2l_{H}^{2}}}{\sqrt{2^{m}m!\sqrt{\pi}l_{H}}}.$

### v2

@@ III.1 Condensate mode
Figure 6: Diagrams for photon scattering from the condensate to a satellite mode with $\omega_{\rm sat}=\omega_{c}$ and nonzero $\bm{q}_{\rm sat}$.
$\varkappa_{\rm sat}$ ${=}\frac{2\pi}{\hbar}\!\!\!\sum_{\begin{subarray}{c}\bm{p}_{n}\\ \bm{q}_{\rm sat}\neq 0\\ {s=\pm 1}\end{subarray}}\!\!\!\left|\!\frac{\lambda_{n,n\shortminus s}^{-}(\bm{q}_{\rm sat})\lambda^{+}_{n\shortminus s,n}(\bm{0})}{E^{s}_{n}{+}s\hbar\omega_{c}{-}E_{n\shortminus s}^{-s}}\right|^{2}$
$\times N^{\rm el}_{n,s}\left(1{-}N^{\rm el}_{n,s}\right)\delta(\hbar\omega_{\rm sat}{-}\hbar\omega_{c}).$ (13)
Here $\bm{q}_{\rm sat}$ denotes the 3D momentum of satellite mode photon.
III.2 The pumping reservoir
Besides being coupled to the condensate mode, the pumping reservoir is exposed to the black-body radiation, leakage through the mirrors and weak photon absorption/emission by the 2DEG (recall that the pumping region is chosen in the ”absorption valley” with low $Q\sim 1$, see Fig. 2). The injection rate from the blackbody source is given by the following expression (here $\bm{v}$ is the photon velocity, $\bm{S}_{\rm res}$ is the oriented area of the sample, $\Delta\omega_{\rm res}$ is the width of the pumping region as shown in Fig. 2):
$W$ ${=}\frac{\Delta N}{\Delta t}{=}\!\!\int\!\!\!\frac{d^{3}p}{(2\pi\hbar)^{3}}\frac{\bm{v}(\bm{p})\cdot\bm{S}_{\rm res}}{\exp\left(\frac{\hbar\omega_{\bm{p}}}{T_{\text{BB}}}\right){-}1}\Theta\!\!\left(\frac{\Delta\omega_{\rm res}}{2}{-}|\omega_{\bm{p}}{-}\omega_{\rm res}|\right)$
$=\varkappa_{\rm res}N_{\rm BE}(T_{\rm BB}),$ (14)
whereas the losses through mirrors for the low Q region are given by
$\varkappa_{\rm res}=\frac{\Delta N}{N_{\rm BE}(T_{\rm BB})\Delta t}$ ${\approx}\frac{\pi\varepsilon cS_{\rm res}\omega_{\rm res}^{2}\Delta\omega_{\rm res}}{(2\pi c)^{3}}.$ (15)
We introduced above $N_{\rm BE}(T_{\rm BB})=(\exp(\hbar\omega_{\rm res}/T_{\rm BB})-1)^{-1}$, representing the thermal occupation of the pumping reservoir that would result from the blackbody acting as the only source.
The photon absorption/emission rates $\varkappa/\varkappa^{\prime}$ are given by the same diagrams and expressions as $\varkappa_{\rm loss}/\varkappa^{\prime}_{\rm loss}$ for the condensate mode (see Fig. 5).
Altogether:
$d_{t}N_{\rm res}$ ${=}W+\varkappa^{\prime}{-}N_{\rm res}\big{(}\varkappa{+}\varkappa_{\rm res}\big{)}$
${+}R(1+N_{0})N_{\rm res}-rN_{0}(1+N_{\rm res}).$ (16)
Here $\varkappa\equiv\varkappa_{\rm loss}(\bm{q}_{\rm res})$ and $\varkappa^{\prime}\equiv\varkappa^{\prime}_{\rm loss}(\bm{q}_{\rm res})$ (see eqs (9)-(III.1)).
@@ III.3 Satellite modes
As mentioned above, when describing condensate coupling to the satellite modes, the coupling rate does not depend on occupations of the latter. That is why we do not consider a kinetic equation for them. In all the further calculations we assume they are thermally populated with occupation $N_{\rm sat}^{\rm th}=N_{\rm BE}(\omega_{c})$.
@@ III.4 Stationary state
$\begin{cases}d_{t}N_{0}\;\;{=}W_{0}+RN_{\rm res}{+}N_{0}\left(N_{\rm res}(R{-}r){-}r{-}\varkappa_{0}-\varkappa_{\rm mirror}\right),\\ d_{t}N_{\rm res}{=}W_{\rm res}{-}N_{\rm res}(R{+}\varkappa{+}\varkappa_{\rm res}){-}N_{0}\left(N_{\rm res}(R{-}r){-}r\right),\end{cases}$ (17)
where we introduced $W_{0}=\varkappa_{{\rm loss}}^{\prime}+\varkappa_{{\rm sat}}N_{\rm sat}^{\rm\rm th}$, $W_{\rm res}=W+\varkappa^{\prime}\simeq W$, $\varkappa_{0}=\varkappa_{\rm loss}+\varkappa_{\rm sat}$ and $\varkappa_{\rm mirror}=\omega_{c}/2\pi Q$.
Stationary occupations of both the condensate and the reservoir are nonzero because of the spontaneous processes and finite due to finite influx $W$. Namely, for small $W=\varkappa_{\rm res}N_{\rm BE}(T_{\rm BB})$ and $W_{0}$, the occupations of both the condensate mode and the reservoir follow the thermal one:
$\begin{split}N_{\rm res}&{=}\frac{\varkappa_{\rm res}(\varkappa_{0}{+}\varkappa_{\rm mirror}{+}r)N_{\rm BE}(T_{\rm BB}){+}W_{0}r}{(\varkappa_{0}+\varkappa_{\rm mirror})R{+}(\varkappa{+}\varkappa_{\rm res})(\varkappa_{0}{+}\varkappa_{\rm mirror}{+}r)},\\ N_{0}&{=}\frac{\varkappa_{\rm res}RN_{\rm BE}(T_{BB}){+}W_{0}(R{+}\varkappa{+}\varkappa_{\rm res})}{(\varkappa_{0}{+}\varkappa_{\rm mirror})R{+}(\varkappa{+}\varkappa_{\rm res})(\varkappa_{0}{+}\varkappa_{\rm mirror}{+}r)}.\end{split}$ (18)
In the opposite limit (for $W\to\infty$), the reservoir occupation is saturated at $N_{\rm max}$
$\begin{split}N_{\rm res}&=N_{\rm max}\equiv\frac{r+\varkappa_{0}+\varkappa_{\rm mirror}}{R-r},\\ N_{0}&=\frac{\varkappa_{\rm res}N_{\rm BE}(T_{\rm BB})-(\varkappa+\varkappa_{\rm res})N_{\rm max}}{\varkappa_{0}+\varkappa_{\rm mirror}}.\end{split}$ (19)
The exact solution of the system (17) for stationary states interpolates between the two regimes as demonstrated in Fig. 7 (c)-(d).
[…]
When identifying the transition point between the two regimes we follow [22] to introduce the lasing and the Bose condensation thresholds $T_{\rm th}^{\rm las/BEC}$ (and the corresponding occupations $N_{0}^{\rm las/BEC}$). The lasing threshold is defined by extrapolating the linear dependence of the condensate occupation on the blackbody temperature $T_{\rm BB}$ as shown on panel (a) in Fig. 7. From (19) one may infer
$N_{BE}(T^{\rm las}_{\rm th})=\left(\frac{\varkappa}{\varkappa_{\rm res}}+1\right)\frac{r+\varkappa_{0}+\varkappa_{\rm mirror}}{R-r}.$ (20)
The BEC threshold is defined by extrapolating the linear dependence of the condensate occupation as a function of the total occupation (of the isolated photonic subsystem: condensate mode, satellite modes and the pumping reservoir) as demonstrated in Fig. 7 (b). One may infer the value $N_{0}^{\rm BEC}$ from the graph and find the corresponding transition temperature, which is defined by
$N_{\rm 0}(T_{\rm th}^{\rm BEC})=N_{0}^{\rm BEC}.$ (21)
@@ IV.1.1 Electron mobility
$\Gamma=\frac{e\hbar}{\mu_{e}m^{*}_{e}}.$ (22)
@@ IV.1.2 Designing a cavity
To create the desired mode configuration, as described in Sec. II, we propose using a heterostructure with high electron mobility inside a Bragg-reflector. This heterostructure hosts several layers of 2DEG (using multiple layers is helpful for increasing overall electron density). A possible pattern of the dielectric constant modulation is depicted in Fig. 8 (a) (half of the structure is shown with a symmetric counterpart implied to the left). In the same plot we show the typical light intensity profile into the properly designed cavity.
@@
[…]
@@ IV.2 Operating regimes
The Fig. 10 shows the regions of condensate existence for various temperatures of the thermal emitter in case of $Q\to\infty$ on the condensate frequency. Note that these regions (in blue) are bounded by both the absorption peaks and the line of threshold pumping level (in red), which is given as a result of solving the equation (see (17))
$N_{\rm res}\!\big{(}R(\omega_{c},\!H){-}r(\omega_{c},\!H)\big{)}{=}\varkappa_{0}(\omega_{c},\!H){+}r(\omega_{c},\!H){+}\varkappa_{\rm mirror}$ (25)
One may clearly see how the area of the condensate phase shrinks with decreasing temperature (note the reversed axis). The critical temperature is $T_{\rm BB}^{\rm crit}=1300$ K (recall that the lattice temperature is $\leavevmode\nobreak\ 300$ K).
Figure 11: The lasing threshold lines for various pumping temperatures. For each specific temperature the lasing state is above the corresponding line.
No real mirror is perfect, which limits the usage of all the regions in Fig. 10. Fig. 11 shows the lasing threshold, where $N_{0}=N_{\rm BE}(T_{\rm th}^{\rm las})$ (see (20)), helping to assess the feasibility of a real light-emitting device for various combinations of control parameters. Notably, neither low nor high magnetic fields are optimal. With decreasing magnetic field the relative width of the absorption minima (see Fig. 2) decreases, thus the losses increase. In contrast, higher magnetic fields shift the pumping region to higher frequencies, where the blackbody spectral density is exponentially suppressed. Therefore, an intermediate optimal magnetic field exists, enabling device operation with a low $Q$ -factor.
One may see how the minimal required blackbody temperature decreases with increasing quality of the mirrors along with a wider range of magnetic fields. For this diagram we considered the largest region in Fig. 10 (since it has the lowest possible operating temperature).
@@ IV.3 Device performance
Now we study the device performance in terms of emission power and efficiency. Since the useful THz radiation is the one that escapes the microcavity through the mirror, the stationary emission intensity $I_{c}=\hbar\omega_{c}\cdot d_{t}n_{0}\big{|}_{\rm mirror}$ is as follows:
$I_{c}=\frac{W\hbar\omega_{c}\varkappa_{\rm mirror}}{\varkappa_{0}+\varkappa_{\rm mirror}}\left(1-\frac{N_{\rm max}}{N_{\rm BE}(T_{\rm BB})}\right).$ (26)
@@
$I_{\rm in}\simeq\hbar\omega_{\rm res}W\left(1-\frac{N_{\rm max}}{N_{\rm BE}(T_{\rm BB})}\right).$ (27)
@@
$\eta=\frac{I_{c}}{I_{\rm in}}\simeq\frac{\omega_{c}}{\omega_{\rm res}}\frac{\varkappa_{\rm mirror}}{\varkappa_{0}+\varkappa_{\rm mirror}}.$ (28)
@@
$Q_{\rm min}=\frac{\omega_{c}}{2\pi(R-r)(N_{\rm BE}(T_{\rm BB})-N_{\rm max})}$ (29)
@@
Operation region of the device is given by thin stripes in the figure 12. These are exactly the valleys between the peaks of condensate photon absorption by the 2DEG as depicted in Fig. 9. These stripes widen with the increase of magnetic field.
@@
Clearly, the practical implementation of a THz-emitting device of the type described in this paper will face multiple challenges. Technical details may strongly affect the final outcome. We outline a few considerations:
  1. 1.
    The threshold $Q$-factor is quite high compared to those achieved experimentally ($\sim 10^{3}$)[40]. However, as described in the same Ref. [40], adding several more layers to the Bragg reflectors may increase $Q$ by orders of magnitude. In addition, one is not constrained by thermal pumping, which we use for demonstrative calculations. Non-equilibrium pumping sources may be used, such as photonic crystal structures that increase the flux on the desired pumping frequency. With increased pumping flux, the required $Q$-factor can be reduced.
  2. 2.
    Another way to increase the thermal input to the cavity is to use THz (metallic) mirrors that reflect the radiation from undesired directions and redirect it into the cavity.
  3. 3.
    In device design, it is desirable to minimize the energy consumed by the lamp to produce radiation at higher frequencies. This may be achieved by placing the lamp inside a cavity made of high-frequency (IR frequency, visible light) Bragg mirrors, which trap the waste radiation. Thus, the power consumption of the lamp decreases once it reaches equilibrium with the cavity.
@@ V Conclusions
In this paper we present a new mechanism for observing microcavity photonic BEC in the THz range. We studied a system of microcavity photons exposed to pumping by a thermal light source. The temperature difference between the cavity and the pumping source drives a photonic flux to the former. Due to interphoton interactions via Landau-quantized electrons in a magnetic field, the system settles down in a stationary state with macroscopic occupation of a single transverse mode of the cavity.
The demonstrated ability to populate a single transverse mode is necessary though insufficient condition for achieving full coherence. In contrast with cold-atom condensate, there is no direct spontaneous symmetry breaking because of negligible interaction between condensate photons. The condensate gain saturation mechanism is purely kinetic (limited influx from the pumping modes), which is closer to the condensation scenario in exciton-polariton condensates.
In this study we treated a set of low-lying modes of the condensate collectively. This definitely precludes the essential discussion on the thermalization thermalization among states with the same transverse momentum and statistics of the condensate. That is why the estimation of the coherence degree of the photonic condensate and investigation of its self-thermalization dynamics are beyond the scope of the current paper. A detailed study of this issue (along the lines of [41] for the photonic condensate in a dye-filled cavity) is a subject of future investigations.
@@ A.1 Electron-phonon
$V^{\perp}_{nm}(k_{z})$ $\sim\int dz\psi^{z}_{m}(z)e^{ik_{z}z/\hbar}\psi_{n}^{z}(z),$
$V^{\parallel}_{nm}(k_{\parallel})$ $\sim\int dy\psi^{y}_{m}(y,p_{x}^{\prime})e^{ik_{y}y/\hbar}\psi_{n}^{y}(y,p_{x})\Big{|}_{p_{x}^{\prime}-p_{x}=k_{x}},$
@@
$\psi^{y}_{m}(y,p_{x})$ $=\frac{H_{m}(\eta)e^{-\eta^{2}}}{\sqrt{2^{m}m!\sqrt{\pi}l_{H}}}.$